OpenStudy (anonymous):

integrate: from 0 to 1 (x-1) / (x+1) = the first step for the solutions says ((x+1) (x+1)) - (2 / (x+1))

OpenStudy (anonymous):

aaaak that is ((x+1)/(x+1))

OpenStudy (anonymous):

\[\int\limits_{0}^{1} (x-1) / (x+1)\] then the next step says: = ((x+1) / (x+1)) - (2/(x+1) = 1 - (2/(x+1) = x-2(ln(x+1).... the only part i get is the last step for integrating the 1 and the 1/x...

OpenStudy (anonymous):

\[\int\limits_{0}^{1} \frac{x-1}{x+1}dx\] \[=\int\limits\limits_{0}^{1} \frac{x-1+(1-1)}{x+1}dx\] You add 1-1 (which is 0) on the top. We added a 0in this way because it doesn't change the expression but it helps us simplify as follows. \[=\int\limits\limits_{0}^{1} \frac{x+1-2}{x+1}dx\] Rearrange the top to look like this Then separate \[=\int\limits\limits\limits_{0}^{1} (\frac{x+1}{x+1}-\frac{2}{x+1})dx\] Now \[=\int\limits\limits\limits_{0}^{1} (1-\frac{2}{x+1})dx\] Do you understand where to go from here?

OpenStudy (anonymous):

ok i thought they were multiplying something not just adding and then subtracting one...that sure seems like a trick question.....thanks

OpenStudy (anonymous):

it does seem like a trick question but as you get to do more and more integral problems, you'll find that this "trick" comes in handy. basically, this is what happened...